(((75=(3.7t^2)/2}}} im not sure what this is supposed to mean...a tutor used it in one of my questions...ive been looking around the sight but it says the expression is unknown... ?
Answer by stanbon(75887)

75=(3.7t^2)/2
3.7t^2 = 150
t^2 = 150/3.7 = 40.54054..
t = +/- 6.367145...
=====================
Cheers,
Stan H.